OpenStudy (tortillabluechip):

Which of the follow is an equivalent function to f(x) = 4(3)^2x? a. f(x) = 4(9)^x b. f(x) = 36^x c. f(x) = 144^x d. f(x) = 4(6x)

OpenStudy (mathmale):

By " f(x) = 4(3)^2x," do you mean f(x) = 4(3)^(2x)? In other words, is your "2x" the exponent of 3, or is only 2 the exponent of 3?